# Modeling of acceleration in heavy-ion collisions: occurrence of temperature below the Unruh temperature

### v1

@@ III.3 Analysis of $T^{2}-T_{U}^{2}$
Since we analyze heavy-ion collisions from the point of view of observing and interpreting the phase transition when $T^{2}<T_{\rm U}^{2}$, the greatest interest is in the analysis of the difference of the squares of the corresponding temperatures $T^{2}-T_{\rm U}^{2}$. We show the profiles of the difference $T^{2}-T_{\rm U}^{2}$ for various $z$-cuts ($z\geq 0$), calculated at various times after the collision of nuclei for both the central ($b=0$ fm, Fig. 3) and off-center ($b=7.5$ fm, Fig. 4) collisions. Let us first consider the central cut ($z=0\,$fm). In the central cut the difference $T^{2}-T_{\rm U}^{2}$ is negative mainly in the outside the fireball while inside the fireball $T^{2}-T_{\rm U}^{2}>0$. There is a very small region of the fireball near the boundary, which disappears by the time $t-t_{coll}=3-4\,$fm$/c$ after the nuclei touch, where $T^{2}-T_{\rm U}^{2}<0$. In the case of an off-center collision, the region inside the fireball, in which $T^{2}-T_{\rm U}^{2}<0$, occupies a somewhat larger relative part of the volume, compared to the case $b=0\,$fm. For an off-center collision at the time $t-t_{coll}=3\,$fm$/c$ inside the fireball regions are still observed in which $T^{2}-T_{\rm U}^{2}<0$. In both the central collision and the off-center collision by the time $t-t_{coll}=5\,$fm$/c$ inside the fireball the difference $T^{2}-T_{\rm U}^{2}$ is positive. Inside the QGP phase region the difference $T^{2}-T_{\rm U}^{2}$ is positive for both the central and off-center collisions.
@@
Figure 3: $T^{2}-T_{\rm U}^{2}$ profiles for different $z$-cuts calculated at various times after the collision for Au + Au central collisions ($b=0\,$fm) at $\sqrt{s_{NN}}=7.7\,$GeV, averaged over all events. Only positive $z$ is shown due to symmetry. The dotted curves on the $T^{2}-T_{\rm U}^{2}$ profiles show two contours in the $xy$ plane, limiting the medium in energy density to the value $\varepsilon=500\,$MeV/fm3 (inner contour) and $\varepsilon=50\,$MeV/fm3 (outer contour).
Figure 4: The same as Fig. 3 but for the non-central collisions ($b=7.5\,$fm).
@@
Let us now discuss the evolution of $T^{2}-T_{\rm U}^{2}$ in the volume of the medium. Due to symmetry, we will consider only cuts with $z>0\,$fm. It can be seen that in the first moments of time after the collision ($t-t_{coll}=0-1\,$fm$/c$) the $T^{2}-T_{\rm U}^{2}$ profiles depend significantly on $z$. The fireball is formed inside the central cell ($z=0\,$fm). In the regions with $z>0\,$fm the medium remains cold and accelerated and does not extend beyond $z=1\,$fm. The QGP phase is concentrated in the central cell ($z=0\,$fm) while the regions with $z>0\,$fm contain hadrons. After $t-t_{coll}=1\,$fm$/c$ the dependence of the difference $T^{2}-T_{\rm U}^{2}$ on $z$ weakens and starting from 222\,2fm$/c$ after the collision the qualitative picture of the difference $T^{2}-T_{\rm U}^{2}$ does not change depending on $z$. The fireball expands and the outer regions remain accelerated. In the presented profiles it can be seen that the transition from the region with $T^{2}-T_{\rm U}^{2}>0$ to the region with $T^{2}-T_{\rm U}^{2}<0$ occurs through the region $T^{2}-T_{\rm U}^{2}=0$. In the central section the latter corresponds to the transition from the region of the QGP phase to the hadron phase at the moment of maximum overlap. Thus, the hot medium containing the QGP lies inside the cold accelerated region consisting of hadrons. The transition between the phases occurs through the region $T^{2}-T_{\rm U}^{2}=0$.
IV Interpretation: phase transition at Unruh temperature
@@
$ds^{2}=\frac{\rho^{2}}{v^{2}}d\phi^{2}+dx^{2}+dy^{2}+d\rho^{2},$ (8)
where the usual (Euclidean) Rindler coordinates $\rho$ and $0<\phi<2\pi$ are introduced, see e.g. [92, 93]. The coordinate $\rho=a^{-1}$, that is, it sets the inverse proper acceleration, and the proper temperature is included into the parameter $\nu=2\pi T/a$. This space is the direct product of a 2-dimensional cone and a 2-dimensional plane $M=R^{2}\otimes C^{2}_{\nu}$.
@@
and $T^{\mu}_{\mu}\neq 0$ now. Moreover, at the point $T=a/2\pi$ the heat capacity undergoes a discontinuity, that is, a second-order phase transition occurs
@@
Such unusual properties of the region $T<a/2\pi$ could cast doubt on how feasible they are in nature. However, the negativity of energy in (10) simply means the predominance of “potential” energy associated with acceleration over “kinetic” energy associated with temperature. And the critical behavior (11) is due to the fact that at the Unruh temperature the lower Matsubara modes become singular at the event horizon of the space (8).
@@
It was suggested [80] that such states are formed in the early stages of the collision of heavy ions, when, due to deceleration of ions, the beams interaction region is characterized by large acceleration, but has not yet had enough time to thermalize. This statement is well confirmed now by the simulation results shown in Figures 1-4: really in the early stages of the collision, the temperature is below the Unruh temperature.
@@
In view of the results obtained, the following question is raised: is it possible to draw conclusions about the existence of a phase transition? Here we also come to interesting observation. As explicitly shown in Figure 2 the region $T<a/2\pi$, at the same time corresponds predominantly to the hadronic phase, and the region $T>a/2\pi$ corresponds to the partonic phase or quark-gluon plasma [^3] Note that there is an interesting correspondence: massive particles, hadrons, lead to a non-zero trace of the stress-energy, just like in phase (10). Thus, we see that there is a close connection between the phase transition (11) and the confinement-deconfinement transition.
@@
This coincidence is not accidental, see discussion in [80]. Indeed, initially after the collision $T<a/2\pi$ and the energy is negative, being concentrated in the hadronic strings. The negative energy levels can be filled, with an increase in the energy of the system (with increasing temperature to Unruh temperature) and the simultaneous formation of particles with a thermal spectrum [95]. Thus, it was suggested that the state $T<a/2\pi$ decays to form a thermal spectrum of hadrons, thus providing rapid thermalization in heavy ion collisions. Since hadronization occurs as a result of the QCD phase transition, two processes (confinement-deconfinement and phase transition (11) at $T=a/2\pi$) occur simultaneously, or, more preferably, transition (11) might provide dual description of the QCD phase transition. The obtained modeling results confirm this statement at least at a qualitative level.
@@ V Conclusions
The spatial distribution of the acceleration has a core-corona structure, and we find states with temperatures below the Unruh temperature. These states consist mostly of hadrons, while states with a temperature above the Unruh temperature contain a partonic phase. At the initial time moment after the collision, the acceleration is maximum in the regions outside the fireball (corona), while the regions of the fireball (core) are practically not accelerated and are hot. Inside the fireball, the temperature $T$ is significantly higher than the Unruh temperature $T_{\rm U}$. This qualitative picture is preserved for collision energies $\sqrt{s_{NN}}=4.5-11.5\,$GeV. However, the value of the characteristic Unruh temperature becomes significantly higher at high collision energies. After 555\,5fm$/c$, when the medium has thermalized, we practically do not observe accelerated regions in the center of the system.
@@ Acknowledgements.
The authors are thankful to Alejandro Ayala for valuable discussions. The calculations has been carried out using the “Govorun” supercomputer provided by the Laboratory of Information Technologies (JINR). The work of N. S. Tsegelnik also was supported in part by the grants for young scientists and specialists of JINR № 25-301-06. The work of G. Yu. Prokhorov, D. A. Shohonov and V. I. Zakharov was supported by Russian Science Foundation Grant No 24-22-00124.

### v2

@@ III.3 Analysis of $T^{2}-T_{U}^{2}$
Since we analyze heavy-ion collisions from the point of view of observing and interpreting the phase transition when $T^{2}<T_{\rm U}^{2}$, the greatest interest is in the analysis of the difference of the squares of the corresponding temperatures $T^{2}-T_{\rm U}^{2}$. We show the profiles of the difference $T^{2}-T_{\rm U}^{2}$ for various $z$-cuts ($z\geq 0$), calculated at various times after the collision of nuclei for both the central ($b=0$ fm, Fig. 3) and off-center ($b=7.5$ fm, Fig. 4) collisions. Let us first consider the central cut (z=0z=0\,fm). In the central cut the difference $T^{2}-T_{\rm U}^{2}$ is negative mainly in the outside the fireball while inside the fireball $T^{2}-T_{\rm U}^{2}>0$. There is a very small region of the fireball near the boundary, which disappears by the time $t-t_{coll}=3-4\,$fm/c/c after the nuclei touch, where $T^{2}-T_{\rm U}^{2}<0$. In the case of an off-center collision, the region inside the fireball, in which $T^{2}-T_{\rm U}^{2}<0$, occupies a somewhat larger relative part of the volume, compared to the case b=0b=0\,fm. For an off-center collision at the time $t-t_{coll}=3\,$fm/c/c inside the fireball regions are still observed in which $T^{2}-T_{\rm U}^{2}<0$. In both the central collision and the off-center collision by the time $t-t_{coll}=5\,$fm/c/c inside the fireball the difference $T^{2}-T_{\rm U}^{2}$ is positive. Inside the QGP phase region the difference $T^{2}-T_{\rm U}^{2}$ is positive for both the central and off-center collisions.
@@
Figure 3: $T^{2}-T_{\rm U}^{2}$ profiles for different $z$-cuts calculated at various times after the collision for Au + Au central collisions (b=0b=0\,fm) at $\sqrt{s_{NN}}=7.7\,$GeV, averaged over all events. Only positive $z$ is shown due to symmetry. The dotted curves on the $T^{2}-T_{\rm U}^{2}$ profiles show two contours in the $xy$ plane, limiting the medium in energy density to the value $\varepsilon=500\,$MeV/fm3 (inner contour) and $\varepsilon=50\,$MeV/fm3 (outer contour).
Figure 4: The same as Fig. 3 but for the non-central collisions (b=7.5b=7.5\,fm).
@@
Let us now discuss the evolution of $T^{2}-T_{\rm U}^{2}$ in the volume of the medium. Due to symmetry, we will consider only cuts with z>0z>0\,fm. It can be seen that in the first moments of time after the collision ($t-t_{coll}=0-1\,$fm/c/c) the $T^{2}-T_{\rm U}^{2}$ profiles depend significantly on $z$. The fireball is formed inside the central cell (z=0z=0\,fm). In the regions with z>0z>0\,fm the medium remains cold and accelerated and does not extend beyond z=1z=1\,fm. The QGP phase is concentrated in the central cell (z=0z=0\,fm) while the regions with z>0z>0\,fm contain hadrons. After $t-t_{coll}=1\,$fm/c/c the dependence of the difference $T^{2}-T_{\rm U}^{2}$ on $z$ weakens and starting from 22\,fm/c/c after the collision the qualitative picture of the difference $T^{2}-T_{\rm U}^{2}$ does not change depending on $z$. The fireball expands and the outer regions remain accelerated. In the presented profiles it can be seen that the transition from the region with $T^{2}-T_{\rm U}^{2}>0$ to the region with $T^{2}-T_{\rm U}^{2}<0$ occurs through the region $T^{2}-T_{\rm U}^{2}=0$. In the central section the latter corresponds to the transition from the region of the QGP phase to the hadron phase at the moment of maximum overlap. Thus, the hot medium containing the QGP lies inside the cold accelerated region consisting of hadrons. The transition between the phases occurs through the region $T^{2}-T_{\rm U}^{2}=0$.
III.4 The direction of acceleration
Let’s also briefly discuss where the acceleration is directed. Typical distribution for the three-dimensional part of a four-dimensional vector $a^{\mu}=(a^{0},\mathbb{a})=(\gamma\dot{\gamma}c,\gamma^{2}\frac{d\mathbb{v}}{dt}+\gamma\dot{\gamma}\mathbb{v})$, where $\gamma=1/\sqrt{1-\mathbb{v}^{2}/c^{2}}$ is the relativistic gamma factor, is shown in the Figure 5. Figure 6 also shows the distribution of different projections. Let’s consider the core and the corona separately.
Figure 5: Three-dimensional distribution of the three-dimensional part of the acceleration vector $a^{\mu}=u^{\nu}\partial_{\nu}u^{\mu}$ for a non-central collision b=7.5b=7.5\,fm at $t-t_{coll}=5\,$fm/$c$.
Figure 6: Various projections of the three-dimensional part of acceleration $a^{\mu}=u^{\nu}\partial_{\nu}u^{\mu}$ for a non-central collision b=7.5b=7.5\,fm at z=5z=5\,fm and $t-t_{coll}=5\,$fm/$c$.
  • Central region. The acceleration of the central dense and hot region is aligned with the direction of the system’s expansion. At small $z$ the transversal component of the acceleration dominates. At larger values of $z$ the $z$-component of the acceleration dominates, increasing to the boundaries of the hot zones. Nonetheless, this acceleration remains at least an order of magnitude lower than that observed in the corona.
  • Corona area. The acceleration in the corona area is mostly opposite directed to the direction of the system’s expansion, especially at values of $z$ close to the maximums. For such zones the $z$-component of the acceleration is an order of magnitude larger than the other projections. However, actually, in the areas most distant from the center, the acceleration structure becomes more chaotic and requires deeper analysis.
It is important to note that the verbal description presented above is significantly simplified. Evidently, the acceleration field demonstrates significant variations in both space and time. Furthermore, at lower collision energies, the acceleration field will be substantially influenced by the presence of spectators. This will result in the additional deceleration along the boundary between the spectators and participants. We plan to study this issue in more detail in the future.
IV Discussion
IV.1 Interpretation: phase transition at Unruh temperature
@@
$ds^{2}=\frac{\rho^{2}}{\nu^{2}}d\varphi^{2}+dx^{2}+dy^{2}+d\rho^{2},$ (8)
where the usual (Euclidean) Rindler coordinates $\rho$ and $0<\varphi<2\pi$ are introduced, see e.g. [94, 95]. The coordinate $\rho=a^{-1}$, that is, it sets the inverse proper acceleration, and the proper temperature is included into the parameter $\nu=2\pi T/a$. This space is the direct product of a 2-dimensional cone and a 2-dimensional plane $M=R^{2}\otimes C^{2}_{\nu}$.
@@
and $T^{\mu}_{\mu}\neq 0$ now. Note that this energy is generally not equivalent to the fluid energy, for example, in equations (5) and (6) and contains the contribution of particles and vacuum, the separation of which is a non-trivial problem. Therefore, the negativeness of the energy density (10) does not contradict the fact that in the PHSD model the energy density is positive.
At the point $T=a/2\pi$ the heat capacity undergoes a discontinuity, that is, a second-order phase transition occurs
@@
Such unusual properties of the region $T<a/2\pi$ could cast doubt on how feasible they are in nature. However, the negativity of energy in (10) simply means the predominance of “potential” energy associated with acceleration over “kinetic” energy associated with temperature. As already mentioned, energy depends on both temperature $T$ and on acceleration $a$. The first of these contributions, “kinteic energy”, includes the usual energy of medium at a certain temperature $T$ (for example, in the case of free massless particles it is given by the Stephen-Boltzmann law). And the second contribution, “potential energy”, is related with effective statistical interactions with acceleration, described by boost generator (for more details see [84, 82]), and can be negative. We are interested in the situation when the second contribution exceeds the first.
[…]
$\phi^{\pm}_{n,s_{1}}\sim e^{i\pi T(2n+1)\tau}J_{\pm[\frac{2\pi T}{a}(n+1/2)-s_{1}/2]}(\xi\rho)\,$, (12)
but only one of these solutions, $\phi^{+}$ or $\phi^{-}$, is finite at the horizon $\rho\to 0$, since Bessel function asymptotically $J_{b}(x)\to x^{b}$ at $x\to 0$. In the formula (12) $s_{1}=\pm 1$ is the eigenvalue of the boost operator, $\tau=\rho\varphi/\nu$ is the imaginary proper time in Rindler coordinates and $\xi$ is defined by momenta and $\lambda$.
The key point is that when passing to the region $T<T_{U}$, the finite solutions $\phi^{+}_{0,1}$ and $\phi^{-}_{-1,-1}$ becomes singular and vice versa. Thus, we change the solution to the modes $\phi^{-}_{0,1}$ and $\phi^{+}_{-1,-1}$ , which leads to a jump-like behavior of the Green function and the mean values. This ultimately leads to the phase transition (11). This situation is repeated for temperatures $T=T_{U}/(2k+1)$, $k=0,1$... . Note that although the described mode behavior was observed in the simplest case of massless free fields, it can be generalized to the massive case, and it is probably a fairly general phenomenon, like the Unruh effect itself. The jump-like behavior of the modes is shown in the Figure 7.
Figure 7: Change in the behavior of $n$-th Matsubara modes when crossing temperatures $T=T_{U}/(2k+1)$ with $k=0,1$.... The significant temperature-dependent part of the solution with $s_{1}=1$ is shown (without exponential factor).
@@
It was suggested [82] that such states with $T<T_{U}$ are formed in the early stages of the collision of heavy ions, when, due to deceleration of ions, the beams interaction region is characterized by large acceleration, but has not yet had enough time to thermalize. This statement is well confirmed now by the simulation results shown in Figures 1-4: really in the early stages of the collision, the temperature is below the Unruh temperature.
@@
In view of the results obtained, the following question is raised: is it possible to draw conclusions about the existence of a phase transition? Here we also come to an interesting observation. As explicitly shown in Figure 2 the region $T<a/2\pi$, at the same time corresponds predominantly to the hadronic phase, and the region $T>a/2\pi$ corresponds to the partonic phase or quark-gluon plasma [^5] Note that there is an interesting correspondence: massive particles, hadrons, lead to a non-zero trace of the stress-energy, just like in phase (10). Thus, we see that there is a close connection between the phase transition (11) and the confinement-deconfinement transition.
@@
This coincidence is not accidental, see discussion in [82]. Indeed, initially after the collision $T<a/2\pi$ and the energy is negative (see (10)), being concentrated in the hadronic strings. The negative energy levels can be filled, with an increase in the energy of the system (with increasing temperature to Unruh temperature) and the simultaneous formation of particles with a thermal spectrum [97]. Thus, it was suggested that the state $T<a/2\pi$ decays to form a thermal spectrum of hadrons, thus providing rapid thermalization in heavy ion collisions. Since hadronization occurs as a result of the QCD phase transition, two processes (confinement-deconfinement and phase transition (11) at $T=a/2\pi$) occur simultaneously, or, more preferably, transition (11) might provide a dual description of the QCD phase transition. The obtained modeling results confirm this statement at least at a qualitative level. We note a recent work [98], where also the relationship between chiral transition and acceleration was studied using the Nambu-Jona-Lasinio model as an example and a conclusion was made about a direct relationship between critical temperature and acceleration.
We note that, according to, the Figures 3 and 4 even at t=5t=5\,fm/$c$ in the central region $T>T_{U}$. Of course, over time, the entire QGP should turn into hadrons. Therefore, it can be assumed that the temperature becomes lower than $T_{U}$ everywhere.
However, we are still considering not so high times $t\sim 5\,$fm/$c$, at which within the framework of the model used the quark-gluon plasma can still be present in the system. Therefore, the fact that in the central region $T>T_{U}$ does not contradict the above-described picture of phase separation. Moreover, Figure 2 confirms the separation of phases with respect to $T_{U}$. However, it should be noted that at the moment we have only some elements of a complementary description of hadronization based on the discussed new phase transition at $T_{U}$ and we do not claim, that hadrons always correspond to $T<T_{U}$ (and vice versa for quark-gluon plasma).
However, generally speaking, the interpretation discussed in this Subsection does not claim to accurately reproduce all the results of the PHSD model; moreover, such a correspondence would be surprising, given the completely different physical foundations. But at the same time we see that it allows us to shed light on some of the results obtained.
IV.2 Origin of the large acceleration in the corona
Thus, as a result of modeling we have shown that a very large acceleration occurs, and, predominantly, the greatest acceleration is achieved in the corona. The natural question is – what is the source of the large acceleration in the corona?
This is a very interesting and non-trivial question. First of all, it should be noted that a similar structure was also observed in the case of vorticity [99] (see also [100]). We have several assumptions about possible sources of large acceleration in the corona:
  • Relativistic effect. Large acceleration in the corona region can be attributed to the fact that, at the boundaries of the system, particles forming the fluids approach speeds close to the speed of light. Consequently, these areas exhibit a significant Lorentz factor, thereby resulting in deceleration caused by a purely relativistic effect.
  • Geometry of the system. Colliding nuclei at the edges of the system are thinner than in the center. Therefore, given that the system stops, it should be expected that the regions associated with the corona should stop in a shorter interaction time than the regions forming the core and therefore experience greater acceleration.
  • Luttinger relation. There is another intriguing possibility: in a state of thermodynamic equilibrium, an effective force arises, associated with the temperature gradient $a\sim\nabla T$, known as the “Luttinger force” [101, 102]. It is very close to the Tolman-Ehrenfest equilibrium criterion [103] in gravitational field, as well as the conditions of global thermodynamic equilibrium [102]. Of course, the system in particle collisions is far from global equilibrium, however, it is possible that Luttinger relation works approximately. It would be interesting if this relation were fulfilled in the core or corona.
Thus, unfortunately, we do not have a definitive final answer to the question of the sources of large acceleration in the corona. However, clarifying this issue is an interesting task for future research.
@@ V Conclusions
The spatial distribution of the acceleration has a core-corona structure, and we find states with temperatures below the Unruh temperature. These states consist mostly of hadrons, while states with a temperature above the Unruh temperature contain a partonic phase. Short after the collision, the acceleration is maximum in the regions outside the fireball (corona), while the regions of the fireball (core) are practically not accelerated and are hot. Inside the fireball, the temperature $T$ is significantly higher than the Unruh temperature $T_{\rm U}$. This qualitative picture is preserved for collision energies $\sqrt{s_{NN}}=4.5-11.5\,$GeV. However, the value of the characteristic Unruh temperature becomes significantly higher at high collision energies. After 55\,fm/c/c, when the medium has thermalized, we practically do not observe accelerated regions in the center of the system.
We also briefly discussed the issue of the direction of acceleration, in particular, in the central region it is predominantly directed in the direction of expansion of the system.
@@ Acknowledgements.
The authors are thankful to Alejandro Ayala, Elena Bratkovskaya, Maxim Chernodub and Evgeni Kolomeitsev for valuable discussions and comments. The calculations has been carried out using the “Govorun” supercomputer provided by the Laboratory of Information Technologies (JINR). The work of N. S. Tsegelnik also was supported in part by the grants for young scientists and specialists of JINR № 25-301-06. The work of G. Yu. Prokhorov, D. A. Shohonov and V. I. Zakharov was supported by Russian Science Foundation Grant No 24-22-00124.